Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image

am1weekly_2001-202_2001-208

Mission Operations Days: 2001/202 to 2001/208
July 21, 2001 00:00:00 GMT to July 27, 2001 00:00:00 GMT


Terra Spacecraft and MODIS Instrument Status:

Terra (AM-1) is in Normal Ops Mode
MODIS is in Science Mode - A Side
ComponentState(s)Comment
BlackbodyA On (290 K); B OffNominal
Calibration ElectronicsA On; B OffNominal
Control ProcessorA On; B OffNominal
Nadir DoorUnlatched, OpenNominal
Space View DoorUnlatched, OpenNominal
Solar Diffuser DoorUnlatched, ClosedNominal
FDDI FormatterA On; B OffNominal
FDDI PortA On; B OffNominal
FIFO Memory#1 & #2 On; #3 & #4 OffNominal
Format ProcessorA On; B OffNominal
Power Supplies#1 On; #2 OffNominal
PV VIS FPAA On; B OffNominal
PV NIR FPAA On; B OffNominal
PV SMIR FPAA On; B OffNominal
PV LWIR FPAA On; B OffNominal
PC LWIR FPAA On; B OffNominal
RC Outgas HeatersAll Off (CS, IS, OS)Nominal
RC LWIR FPA HeaterOn (83 K)Nominal
RC SMIR FPA HeaterOffNominal
Scan AssemblyA On; B OffNominal
SDSMA Off; B OffNominal
SRCAA Off; B OffNominal
PS1 Survival HeaterEnabledNominal
PS2 Survival HeaterEnabledNominal
Timing GeneratorA Off; B OnNominal
Flight SoftwareRev BD + 1 set of patches
  • FR Error Counting and Statistics (2 patches)
Nominal
Inhibit Ids SetNoneNominal
TMON 66 (SRCA A)DisabledNominal
TMON 67 (SRCA B)DisabledNominal
TMON 68 (FIFO)EnabledNominal
TMON 69 (EPS Current)DisabledNominal

This Week's Completed MODIS Activities:

Saturday, July 21, 2001

None

Sunday, July 22, 2001

None

Monday, July 23, 2001

None

Tuesday, July 24, 2001

None

Wednesday, July 25, 2001

2001/206 18:09:12-18:20:40 ATC OA-15: SD/SDSM Open
2001/206 19:48:06-19:59:34 ATC OA-16: SD/SDSM Screened

Thursday, July 26, 2001

None

Friday, July 27, 2001

An additional 20+ minutes for FR Day Rate was collected during S/C Ground Track Night.
2001/208 14:00:09-14:28:00 ATC FR Night Rate (21:38 minutes less than normal)
2001/208 14:28:00-15:39:02 ATC FR Day Rate (21:38 minutes more than normal)

This Week's Scheduled MODIS Activities Not Completed:

None

Upcoming MODIS Events:

During the upcoming week, it is planned to collect an extra 20 minutes of DAY rate data while the spacecraft is over the night side of the Earth, using SMIR Itwk/Vdet of 79/110 (to be restored to current 79/190 values after data collection). Solar Diffuser calibrations will likely occur on Wednesday, although the precise times have not yet been determined.

Additionally, the Band5 Gains will be decreased by 25% (returning them to the values they were at prior to the PS2 Shutdown Anomaly). This will be done as time permits on either Monday or Tuesday.

Friday, July 27, 2001

None

Saturday, July 28, 2001

None

Sunday, July 29, 2001

None

Monday, July 30, 2001

None

Tuesday, July 31, 2001

None

Wednesday, August 1, 2001

None

Thursday, August 2, 2001

None

Friday, August 3, 2001

None

Maneuvers:

The next MODIS Lunar Roll Calibration is scheduled for August 9, 2001 ( DOY 221 ) from 05:24:35 to 05:31:05. A second lunar calibration will be performed without a roll maneuver on the same day from 20:23:37 to 20:30:07. During the second roll, the SMIR Itwk/Vdet will be temporarily set to 70/110.

MODIS Anomalies:

Investigation into the PS2 shutdown anomaly is still underway. An anomaly report is expected by the end of this month.

General Instrument Comments:

MODIS is in Science Mode on the A-side with the SVD and NAD open, operating with the Formatter Resolution Flight Software Patches (FRFSP).

MODIS Telemetry Trends:

Trending has resumed with the instrument on the A-side. All appears nominal at this time.

Non-MODIS Significant Events:

None.

Limited Life Item Status:

SRCA 10W Lamp #1: 189.0 of 500 hours
SRCA 10W Lamp #2: 144.7 of 500 hours
SRCA 10W Lamp #3: 153.9 of 500 hours
SRCA 10W Lamp #4: 61.5 of 500 hours

SRCA 1W Lamp #1: 557.9 of 4000 hours
SRCA 1W Lamp #2: 276.3 of 4000 hours

Solar Diffuser Door: 1657 of 3022 Movements
Nadir Aperture Door: 532 of 1316 Movements
Space View Door: 437 of 1316 Movements